# Pseudo-Boolean Proof Logging for Optimal Classical Planning

### v1

@@ Introduction
For the correctness of plans, such a verification is performed routinely: the generated plans themselves serve as certificates for this, and plan validation tools such as VAL (Howey and Long 2003) or INVAL (Haslum 2017) can be used to check that the produced plans are correct. For unsolvability, Eriksson et al. (Eriksson, Röger, and Helmert 2017, 2018; Eriksson and Helmert 2020) recently introduced two forms of unsolvability certificates that cover very diverse planning algorithms.
@@
The certificate is based on cutting planes proofs with reification (Bogaerts et al. 2023), in which the atomic pieces of knowledge are pseudo-Boolean constraints. Such constraints allow us to reason conveniently about costs of actions and bounds on costs to reach a state. Roughly speaking, the certificate describes an overapproximation of which states can be reached at which cost and shows that no goal state can be reached at a cost below the claimed optimal plan cost. The heart of the proof is a standard pseudo-Boolean proof that can be verified by an off-the-shelf proof checker such as the formally verified VeriPB system.
@@
A Boolean variable has domain $\{0,1\}$. A literal is a Boolean variable $x$ or its negation $\bar{x}$. A pseudo-Boolean (PB) constraint (in normalized form) over a finite set $X=\{x_{1},\dots,x_{n}\}$ of Boolean variables is an inequality
@@
To summarize, a CPR proof consists of a sequence of derivation steps from the cutting plane proof system (literal axiom, linear combination, division, and saturation), constraints derived by RUP, and reifications. In addition, we allow the redundance-based strengthening (RED) rule from VeriPB, which can be understood as a form of proof by contradiction. We only use it in the supplemental material and therefore describe it there.
@@
The PB encoding of task $\Pi=\langle\mathcal{V},\mathcal{A},I,G\rangle$ uses the propositional variables from $\mathcal{V}$ as Boolean variables, and variables $\mathcal{V}_{c}=\{c_{0},\dots,c_{\lceil\log_{2}(B)\rceil}\}$ as a binary representation of a number in the range $0,\dots,B$. These variables allow us to represent pairs $\langle s,c\rangle$ consisting of a state $s$ and a number $c\leq B$, with the intuition that whenever such a pair is represented by a model in our encoding, we consider it possible that state $s$ is reachable from the initial state at cost $c$ while staying below the bound $B$. In addition, we introduce a number of reification variables.
@@
A certificate shows that the task is unsolvable under a cost bound $B$, i.e., there is no plan $\pi$ with $\textit{cost}(\pi)<B$. Intuitively, it encodes a set of pairs $\langle s,c\rangle$ of states and costs such that (1) it covers the initial state with cost 00, (2) it covers no goal state with a cost strictly lower than $B$, and (3) whenever $\langle s,c\rangle$ is covered and action $a$ is applicable in $s$, then $s\llbracket a\rrbracket$ is covered with cost $c+\textit{cost}(a)$ if this is below $B$. The certificate must prove these three properties by means of three separate CPR proofs. We call property (1) the initial state lemma, property (2) the goal lemma and property (3) the inductivity lemma.
The set of pairs $\langle s,c\rangle$ is defined by a sequence of reifications. Alternatively, we can think of it as a circuit where each gate evaluates a PB constraint. The state variables and cost bits $c_{i}$ are the inputs to the circuit, and the output of the circuit determines whether the input state-cost pair is contained in the set or not. We now formalize this notion.
@@ A PB circuit with input variables $V$ is a pair $\mathcal{R}=\langle R,r\rangle$, where
Slightly abusing notation, we also interpret a sequence of PB constraints as the PB formula consisting of its components (i.e., treat the sequence as a set). For a PB circuit $\langle R,r\rangle$ with input variables $\mathcal{V}\cup\mathcal{V}_{c}$, we write $M(R,r)$ for the represented set of pairs $\langle s,c\rangle$. In other words, for each model $\rho$ of $R\cup\{r=1\}$, the set $M(R,r)$ contains the pair $\langle s,c\rangle$ where $s=\{v\in\mathcal{V}\mid\rho(v)=1\}$ and $c=\sum_{i}2^{i}\rho(c_{i})$.
@@
    $\mathcal{C}_{\textup{init}}\cup\mathcal{C}_{\varphi}\models(r_{I}\land\overline{\textit{cost}_{\geq 1}})\rightarrow r_{\varphi}$
@@
    $\mathcal{C}_{\textup{goal}}\cup\mathcal{C}_{\varphi}\models(r_{G}\land\overline{\textit{cost}_{\geq B}})\rightarrow\overline{r_{\varphi}}$
@@
    $\mathcal{C}_{\textup{trans}}\cup\mathcal{C}_{\varphi}\cup\mathcal{C}_{\varphi}^{\prime}\models(r_{\varphi}\land r_{T})\rightarrow r_{\varphi}^{\prime}$, where $\mathcal{C}_{\varphi}^{\prime}$ is a copy of $\mathcal{C}_{\varphi}$ where all PB variables are replaced with their primed version.
@@
Any solution for the task starts from the initial state with cost 00, which is covered by the certificate according to the initial state lemma. The PB encoding of the task is such that exactly the action applications that do not exceed the cost bound are captured by $r_{T}$. The inductivity lemma establishes that the resulting state-cost pair will always be covered by the certificate, so the certificate covers all states that are reachable below the cost bound. The goal lemma establishes that there is no goal state among these states, so there is no solution of cost strictly below $B$. ∎
@@
The following two lemmas establish that these clipped costs behave as expected. [^1] Throughout the paper, we omit proofs of technical lemmas that do not provide further insight. These proofs are included in the supplemental material. Intuitively, the first lemma says that if the costs represented by the variables in $\mathcal{V}_{c}$ exceed $j+k$, they also exceed the smaller value $j$:
@@
Parts of the overall proof must be contributed by the heuristic. Throughout its execution (a potential initialization phase and evaluations for a number of states), the heuristic maintains its own PB circuit $\langle H,r^{h}\rangle$ with input variables $\mathcal{V}$ and $\mathcal{K}$. We assume that the heuristic uses its own namespace, so it does not introduce reification variables that are also introduced by the search or some other heuristic. Whenever the search uses the heuristic to evaluate a state $s$, the heuristic not only returns the estimate but also a PB variable $r^{h}_{s}$ with the following requirements:
@@
    $\mathcal{C}_{\textup{init}}\cup H\cup\mathcal{C}_{K}\cup\mathcal{C}_{s}\models(r_{s}\land\textit{cost}_{\geq\max\{0,B-h(s)\}})\rightarrow r_{s}^{h}$, where $\mathcal{C}_{s}=\{r_{s}\Leftrightarrow\sum_{v\in s}v+\sum_{v\in\mathcal{V}\setminus s}\bar{v}\geq|\mathcal{V}|\}$.
@@
    $\mathcal{C}_{\textup{goal}}\cup H\cup\mathcal{C}_{K}\models(r_{G}\land\overline{\textit{cost}_{\geq B}})\rightarrow\overline{r_{s}^{h}}$.
@@
    $\mathcal{C}_{\textup{trans}}\cup H\cup H^{\prime}\cup\mathcal{C}_{K}\cup\models(r_{s}^{h}\land r_{T})\rightarrow{r^{\prime}}_{s}^{h}$,
@@
It is not a coincidence that these heuristic certificates resemble lower-bound certificates under cost bound $B$: they can be seen as such certificates for state $s$, where some cost of at least $B-h(s)$ has already been spent to reach $s$.
@@
to $A$, characterizing all pairs $\langle s,\tilde{g}\rangle$ with $\tilde{g}\geq g$. In addition, it adds $\langle s,g\rangle$ to the initially empty collection Closed. It also logs some derivations in $L$ that we describe later. If a successor $s^{\prime}$ with g-value $g^{\prime}$ is not added to open because $\textup{A}^{*}$’s duplicate detection is aware of an earlier encounter with $g^{\prime\prime}<g^{\prime}$, it uses Lemma Proof-Logging Heuristic Search and logs
@@
In the following we explain how we can generate the three proofs $\mathcal{P}_{\textup{init}},\mathcal{P}_{\textup{ind}}$ and $\mathcal{P}_{\textup{goal}}$ for the PB circuit $\langle A,r_{\textup{A}^{*}}\rangle$. We start by showing how $L$ can be extended to $\mathcal{P}_{\textup{init}}$.
@@
RUP can derive the initial state lemma $(r_{I}\land\overline{\textit{cost}_{\geq 1}})\rightarrow r_{\textup{A}^{*}}$ from $\mathcal{C}_{\textup{init}}\cup A$.
@@
Assume (a) $r_{I}\geq 1$, (b) $\overline{\textit{cost}_{\geq 1}}\geq 1$, and (c) $\overline{r_{\textup{A}^{*}}}\geq 1$. From (a) and (1), we receive (d) $v\geq 1$ for each $v\in I$ and $\overline{v}\geq 1$ for each $v\in\mathcal{V}\setminus I$. Reification (4) and (b) express $\sum\nolimits_{i=0}^{\lceil\log_{2}B\rceil}2^{i}c_{i}<1$, from them we get for all $c_{i}\in\mathcal{V}_{c}$ that $\overline{c_{i}}\geq 1$. With (4), we receive $\textit{cost}_{\geq 0}$, which with (d) and (11) gives $r_{I,\textup{g}\geq 0}\geq 0$. Since $\langle I,0\rangle$ is in Closed, this yields with (a) and (13) $r_{\textup{A}^{*}}\geq 1$, contradicting (c). ∎
@@
It is possible to derive the goal lemma $(r_{G}\land r_{\textup{A}^{*}})\rightarrow\textit{cost}_{\geq B}$ from $\mathcal{C}_{\textup{goal}}\cup A$.
@@
For every action $a$ and state $s$ closed with cost $g$, it is possible to derive $(r_{s,\textup{g}\geq g}\land r_{a})\rightarrow r^{\prime}_{\textup{A}^{*}}$ from $\mathcal{C}_{\textup{trans}}\cup A$.
@@
From (b), we can derive with (11) that (e) $v\geq 1$ for all $v\in s$, (f) $\overline{v}\geq 1$ for all $v\in\mathcal{V}\setminus s$, and (g) $\textit{cost}_{\geq g}\geq 1$. We use (c) with (7) to derive (h) $\Delta c^{=\textit{cost}(a)}\geq 1$. With (h), (g), and (a) we get (i) $\textit{cost}^{\prime}_{\geq\min\{B,g+\textit{cost}(a)\}}\geq 1$. If $s$ is a goal state then $g=B$ and we get from (i) and (7) that $\overline{r_{a}}\geq 1$, contradicting (c). Otherwise proceed as follows:
If $a$ is not applicable in $s$, some precondition is violated and we use (f) and (7) to derive $\overline{r_{a}}\geq 1$, contradicting (c).
Otherwise, $a$ is applicable. We define $\tilde{s}=s\llbracket a\rrbracket$ and use (7), (c), (e), (f), (g), and (6) to derive that (j) the primed state variables encode $\tilde{s}$ and that (k) $K^{\prime}_{\geq g+\textit{cost}(a)}\geq 1$.
@@
For every state $s$ closed with cost $g$, it is possible to derive $(r_{s,\textup{g}\geq g}\land r_{T})\rightarrow r^{\prime}_{\textup{A}^{*}}$ from $\mathcal{C}_{\textup{trans}}\cup A$.
@@
We establish by Lemma 3 that (a) $(r_{s,\textup{g}\geq g}\land r_{a})\rightarrow r^{\prime}_{\textup{A}^{*}}$ for every action $a\in\mathcal{A}$.
Then RUP derives the constraint: Assume (b) $r_{s,\textup{g}\geq g}\geq 1$, (c) $r_{T}\geq 1$, and (d) $\overline{r^{\prime}_{\textup{A}^{*}}}\geq 1$. With (a), (b) and (d) we derive $\overline{r_{a}}\geq 1$ for each action $a$. With (8), this yields $\overline{r_{T}}\geq 1$, contradicting (c). ∎
@@
It is possible to derive the inductivity lemma $(r_{\textup{A}^{*}}\land r_{T})\rightarrow r^{\prime}_{\textup{A}^{*}}$ from $\mathcal{C}_{\textup{trans}}\cup A$.
@@
In the proof of the inductivity lemma we first use as many constant-size statements as there are open states. The second RUP subproof in the inductivity lemma requires the constraint from Lemma 4 specified on each state $s\in\textit{Closed}$, which in turn requires the constraint from Lemma 3 specified on each applicable action in $s$. This amortizes with the generation of the successors of $s$ when it moves from Open to Closed. We see that in total there is a constant-factor overhead to the proof-logging of $\textup{A}^{*}$.

### v2

@@ Introduction
For the correctness of plans, such a verification is performed routinely: the generated plans themselves serve as certificates for this, and plan validation tools such as VAL (Howey and Long 2003) or INVAL (Haslum 2016) can be used to check that the produced plans are correct. For unsolvability, Eriksson et al. (Eriksson, Röger, and Helmert 2017, 2018; Eriksson and Helmert 2020) recently introduced two forms of unsolvability certificates that cover very diverse planning algorithms.
@@
The certificate is based on cutting planes proofs with reification (Bogaerts et al. 2023), in which the atomic pieces of knowledge are pseudo-Boolean constraints. Such constraints allow us to reason conveniently about costs of actions and bounds on costs to reach a state. Roughly speaking, the certificate describes an overapproximation of which states can be reached at which cost and shows that no goal state can be reached at a cost below the claimed optimal plan cost. The heart of the proof is a standard pseudo-Boolean proof that can be verified by an off-the-shelf proof checker such as the VeriPB system.
The paper is structured as follows: we introduce the general framework for certifying plan optimality by providing an encoding of planning tasks as pseudo-Boolean constraints and defining lower-bound certificates. Next, we define heuristic certificates, which represent the concept of an admissible heuristic estimate as pseudo-Boolean constraints. We then show how heuristic certificates can be used to integrate proof-logging into $\textup{A}^{*}$, capturing the arguments for optimality of $\textup{A}^{*}$ with an admissible heuristic on a given state space as pseudo-Boolean constraints. Finally, we describe how heuristic certificates for pattern database heuristics and $h^{\textup{max}}$ can be generated, which amounts to proving the admissibility of these heuristics with pseudo-Boolean constraints.
@@
A Boolean variable has domain $\{0,1\}$. A literal $\ell$ is a Boolean variable $x$ or its negation $\bar{x}$. The negation of a literal $\ell=\bar{x}$ is ¯$\bar{\ell}=x$. A pseudo-Boolean (PB) constraint (in normalized form) over a finite set $X=\{x_{1},\dots,x_{n}\}$ of Boolean variables is an inequality
@@
To summarize, a CPR proof consists of a sequence of derivation steps from the cutting plane proof system (literal axiom, linear combination, division, and saturation), constraints derived by RUP, and reifications. In addition, we allow the redundance-based strengthening (RED) rule from VeriPB, which can be understood as a form of proof by contradiction. We only use it in the appendix and therefore describe it there.
@@
The PB encoding of task $\Pi=\langle\mathcal{V},\mathcal{A},I,G\rangle$ uses the propositional variables from $\mathcal{V}$ as Boolean variables, and variables $\mathcal{V}_{c}=\{c_{0},\dots,c_{\lceil\log_{2}(B)\rceil}\}$ as a binary representation of a number in the range $0,\dots,B$. These variables allow us to represent pairs $\langle s,c\rangle$ consisting of a state $s$ and a number $c\leq B$. In the following, one can use the intuition that $s$ is a state of the task, which has been reached incurring cost $c$. In addition, we introduce a number of reification variables.
@@
A certificate shows that the task is unsolvable under a cost bound $B$, i.e., there is no plan $\pi$ with $\textit{cost}(\pi)<B$.
Intuitively, it is based on an invariant $\varphi$ which represents an overapproximation of the reachable state-cost pairs. An invariant in general is a property that is preserved through action applications, i.e. (1) whenever $\varphi$ is true for state $s$ and cost $c$, and action $a$ is applicable in $s$, then $\varphi$ is true for state $s\llbracket a\rrbracket$ and cost $c+\textit{cost}(a)$. [^1] For unsolvability under a cost bound, it is sufficient to only require the invariant property for action applications within the cost bound. We implicitly do this because our task encoding only permits such action applications, cf. Eq. (7). In addition, we require that (2) $\varphi$ is true for the initial state and cost 00. Together, (1) and (2) ensure that $\varphi$ is true in all reachable state-cost pairs. If in addition, (3) $\varphi$ is not true for any goal state with a cost strictly lower than $B$, this implies that the task cannot be solved with cost $<B$: initially, the invariant is true (2), it is impossible to change this by an action application (1), so for all reachable goal states the incurred cost is at least $B$ (3).
In our certificates, the invariant will be defined by a sequence of PB reifications, and the certificate must prove the three properties above by means of three separate CPR proofs. We call property (1) the inductivity lemma, property (2) the initial state lemma and property (3) the goal lemma.
We can think of the representation of the invariant as a circuit where each gate evaluates a PB constraint. The state variables and cost bits $c_{i}$ are the inputs to the circuit, and the output of the circuit determines whether the invariant is true for the input state-cost pair. We now formalize this notion.
@@ A PB circuit with input variables $V$ is a pair $\mathcal{R}=\langle R,r\rangle$, where
Slightly abusing notation, we also interpret a sequence of PB constraints as the PB formula consisting of its components (i.e., treat the sequence as a set). For a PB circuit $\langle R,r\rangle$ with input variables $\mathcal{V}\cup\mathcal{V}_{c}$, we will write $M(R,r)$ for the represented set of pairs $\langle s,c\rangle$. In other words, $M(R,r)$ contains the pair $\langle s,c\rangle$ iff there exists a model $\rho$ of $R\cup\{r=1\}$, such that $s=\{v\in\mathcal{V}\mid\rho(v)=1\}$ and $c=\sum_{i}2^{i}\rho(c_{i})$.
Putting the pieces together, we get the following formal definition of lower-bound certificates. Remember that in the encoding of the planning task, $\mathcal{V}$ contains the state variables and $\mathcal{V}_{c}$ the binary variables for representing cost.
@@
    $\mathcal{C}_{\textup{init}}\cup\mathcal{C}_{\varphi}\cup\mathcal{C}_{\geq}\models(r_{I}\land\overline{\textit{cost}_{\geq 1}})\rightarrow r_{\varphi}$.
@@
    $\mathcal{C}_{\textup{goal}}\cup\mathcal{C}_{\varphi}\cup\mathcal{C}_{\geq}\models(r_{G}\land{r_{\varphi}})\rightarrow{\textit{cost}_{\geq B}}$.
@@
    $\mathcal{C}_{\textup{trans}}\cup\mathcal{C}_{\varphi}\cup\mathcal{C}_{\varphi}^{\prime}\cup\mathcal{C}_{\geq}\models(r_{\varphi}\land r_{T})\rightarrow r_{\varphi}^{\prime}$, where $\mathcal{C}_{\varphi}^{\prime}$ is a copy of $\mathcal{C}_{\varphi}$ where all PB variables are replaced with their primed version.
@@
Assume there is a plan $\pi=\langle a_{1},\dots,a_{n}\rangle$ with $\textit{cost}(\pi)<B$. It induces a sequence of state-cost pairs $\langle s_{0},g_{0}\rangle,\dots,\langle s_{n},g_{n}\rangle$ such that $s_{0}$ is the initial state, $g_{0}=0$, and for all $i\in\{1,\dots,n\}$, $s_{i}=s_{i-1}\llbracket a_{i}\rrbracket$ and $g_{i}=g_{i-1}+\textit{cost}(a_{i})$. Note that $s_{n}$ is a goal state and $g_{n}=\textit{cost}(\pi)<B$.
By the initial state lemma, $\langle s_{0},g_{0}\rangle$ is in $M(\mathcal{C}_{\varphi},r_{\varphi})$. In $r_{T}$, the PB encoding captures all action applications that do not exceed the cost bound, which is the case for the action applications in $\pi$ because $\textit{cost}(\pi)<B$. Thus, by the inductivity lemma, we get that $\langle s_{i},g_{i}\rangle\in M(\mathcal{C}_{\varphi},r_{\varphi})$ for all $i\in\{1,\dots,n\}$. By the goal lemma, we have for all $\langle s,g\rangle\in M(\mathcal{C}_{\varphi},r_{\varphi})$ that $g\geq B$ if $s$ is a goal state. Since $s_{n}$ is a goal state, and $\langle s_{n},g_{n}\rangle\in M(\mathcal{C}_{\varphi},r_{\varphi})$, this implies that $g_{n}\geq B$, a contradiction to $g_{n}<B$. ∎
@@
The following two lemmas establish that these clipped costs behave as expected. [^2] Throughout the paper, we omit proofs of technical lemmas that do not provide further insight. These proofs are included in the appendix. Intuitively, the first lemma says that if the costs represented by the variables in $\mathcal{V}_{c}$ exceed $j+k$, they also exceed the smaller value $j$:
@@
Parts of the overall proof must be contributed by the heuristic. Intuitively, this contribution can be seen as a certificate that shows that it is impossible to reach the goal with a total cost strictly less than $B$ from a state $s$ that has been reached incurring a cost of at least $c$. For an admissible heuristic, this will naturally be the case for all $c$ where $c+h(s)\geq B$. We will later see examples for different heuristics where some use the same invariant for all states and others generate a new invariant for each evaluated state.
Throughout its execution (a potential initialization phase and evaluations for a number of states), the heuristic maintains its own PB circuit $\langle H,r^{h}\rangle$ with input variables $\mathcal{V}$ and $\mathcal{K}$. We assume that the heuristic uses its own namespace, so it does not introduce reification variables that are also introduced by the search or some other heuristic. Whenever the search uses the heuristic to evaluate a state $s$, the heuristic not only returns the estimate but also a PB variable $r^{h}_{s}$ with the following requirements:
@@
    $\mathcal{C}_{s}\cup H\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}\models(r_{s}\land\textit{cost}_{\geq\max\{0,B-h(s)\}})\rightarrow r_{s}^{h}$, where $\mathcal{C}_{s}=\{r_{s}\Leftrightarrow\sum_{v\in s}v+\sum_{v\in\mathcal{V}\setminus s}\bar{v}\geq|\mathcal{V}|\}$.
@@
    $\mathcal{C}_{\textup{goal}}\cup H\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}\models(r_{G}\land{r_{s}^{h}})\rightarrow{\textit{cost}_{\geq B}}$.
@@
    $\mathcal{C}_{\textup{trans}}\cup H\cup H^{\prime}\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}\models(r_{s}^{h}\land r_{T})\rightarrow{r^{\prime}}_{s}^{h}$,
@@
As mentioned earlier, heuristic certificates can be seen as certificates for state $s$ where some cost of at least $B-h(s)$ has already been spent to reach $s$. Thus it is no coincidence that these heuristic certificates also structurally resemble lower-bound certificates under cost bound $B$.
@@
The invariant from a proof-logging $\textup{A}^{*}$ search will conceptually cover two aspects: (1) the invariant is true for all closed states with their corresponding cost from the initial state. For these states, the heuristic estimates cannot rule out that they could be traversed with the corresponding cost by an optimal plan. (2) the invariant is also true for any state-cost pair for which any invariant is true that was produced by the heuristic to actually rule states out. The first part considers the distances from the initial state to the states in the closed list, while the second part considers the distances from the states in the open list to a goal.
@@
to $A$, characterizing all pairs $\langle s,\tilde{g}\rangle$ with $\tilde{g}\geq g$. In addition, it adds $\langle s,g\rangle$ to the initially empty collection Closed. It also logs some derivations in $L$ that we describe later. If a successor $s^{\prime}$ with g-value $g^{\prime}$ is not added to Open because $\textup{A}^{*}$’s duplicate detection is aware of an earlier encounter with $g^{\prime\prime}\leq g^{\prime}$, it uses Lemma Proof-Logging Heuristic Search and logs
@@
In the following we explain how we can generate the three proofs $\mathcal{P}_{\textup{init}},\mathcal{P}_{\textup{ind}}$ and $\mathcal{P}_{\textup{goal}}$ for the PB circuit $\langle A,r_{\textup{A}^{*}}\rangle$. This generation relies on the correctness of $\textup{A}^{*}$. However, the verifier receiving the generated proofs relies only on the correctness of the derivation rules of CPR. We start by showing how $L$ can be extended to $\mathcal{P}_{\textup{init}}$.
@@
RUP can derive the initial state lemma $(r_{I}\land\overline{\textit{cost}_{\geq 1}})\rightarrow r_{\textup{A}^{*}}$ from $\mathcal{C}_{\textup{init}}\cup A\cup\mathcal{C}_{\geq}$.
@@
Assume (a) $r_{I}\geq 1$, (b) $\overline{\textit{cost}_{\geq 1}}\geq 1$, and (c) $\overline{r_{\textup{A}^{*}}}\geq 1$. From (a) and (1), we receive (d) $v\geq 1$ for each $v\in I$ and $\overline{v}\geq 1$ for each $v\in\mathcal{V}\setminus I$. Reification (4) and (b) express $\sum\nolimits_{i=0}^{\lceil\log_{2}B\rceil}2^{i}c_{i}<1$, from them we get for all $c_{i}\in\mathcal{V}_{c}$ that $\overline{c_{i}}\geq 1$. With (4), we receive $\textit{cost}_{\geq 0}$, which with (d) and (11) gives $r_{I,\textup{g}\geq 0}\geq 0$. Since $\langle I,0\rangle$ is in Closed, with (a) and (13) we get $r_{\textup{A}^{*}}\geq 1$, contradicting (c). ∎
@@
It is possible to derive the goal lemma $(r_{G}\land r_{\textup{A}^{*}})\rightarrow\textit{cost}_{\geq B}$ from $\mathcal{C}_{\textup{goal}}\cup A\cup\mathcal{C}_{\geq}$.
@@
For every action $a$ applicable in $s$ and state $s$ closed with cost $g$, it is possible to derive $(r_{s,\textup{g}\geq g}\land r_{a})\rightarrow r^{\prime}_{\textup{A}^{*}}$ from $\mathcal{C}_{\textup{trans}}\cup A\cup A^{\prime}\cup\mathcal{C}_{\geq}$.
@@
From (b), we can derive with (11) that (e) $v\geq 1$ for all $v\in s$, (f) $\overline{v}\geq 1$ for all $v\in\mathcal{V}\setminus s$, and (g) $\textit{cost}_{\geq g}\geq 1$. We use (c) with (7) to derive (h) $\Delta c^{=\textit{cost}(a)}\geq 1$. With (h), (g), and (a) we get (i) $\textit{cost}^{\prime}_{\geq\min\{B,g+\textit{cost}(a)\}}\geq 1$. If $s$ is a goal state then $g=B$ and we get from (i) and (7) that $\overline{r_{a}}\geq 1$, contradicting (c).
Otherwise, we define $\tilde{s}=s\llbracket a\rrbracket$ and use (7), (c), (e), (f), (g), and (6) to derive that (j) the primed state variables encode $\tilde{s}$ and that (k) $K^{\prime}_{\geq g+\textit{cost}(a)}\geq 1$.
@@
For every state $s$ closed with cost $g$, it is possible to derive $(r_{s,\textup{g}\geq g}\land r_{T})\rightarrow r^{\prime}_{\textup{A}^{*}}$ from $\mathcal{C}_{\textup{trans}}\cup A\cup A^{\prime}\cup\mathcal{C}_{\geq}$.
@@
We establish by Lemma 3 that (a) $(r_{s,\textup{g}\geq g}\land r_{a})\rightarrow r^{\prime}_{\textup{A}^{*}}$ for every action $a\in\mathcal{A}$ that is applicable in $s$.
Then RUP derives the constraint: Assume (b) $r_{s,\textup{g}\geq g}\geq 1$, (c) $r_{T}\geq 1$, and (d) $\overline{r^{\prime}_{\textup{A}^{*}}}\geq 1$. With (a), (b) and (d) we derive $\overline{r_{a}}\geq 1$ for each action $a$ applicable in $s$. From (b), we can derive with (11) that (e) $v\geq 1$ for all $v\in s$, (f) $\overline{v}\geq 1$ for all $v\in\mathcal{V}\setminus s$, and if $a$ is not applicable in $s$, some precondition is violated and we use (f) and (7) to derive $\overline{r_{a}}\geq 1$. With (8), this yields $\overline{r_{T}}\geq 1$, contradicting (c). ∎
@@
It is possible to derive the inductivity lemma $(r_{\textup{A}^{*}}\land r_{T})\rightarrow r^{\prime}_{\textup{A}^{*}}$ from $\mathcal{C}_{\textup{trans}}\cup A\cup A^{\prime}\cup\mathcal{C}_{\geq}$.
@@
In the proof of the inductivity lemma we first use as many constant-size statements as there are open states. The second RUP subproof in the inductivity lemma requires the constraint from Lemma 4 specified on each state $s\in\textit{Closed}$, which in turn requires the constraint from Lemma 3 specified on each applicable action in $s$. This amortizes with the generation of the successors of $s$ when $s$ moves from Open to Closed. We see that in total there is a constant-factor overhead to the proof-logging of $\textup{A}^{*}$.